OpenStudy (anonymous):

In right triangle ABC, C is the right angle. BC=5 and AB=13. What is the value of sec A?

OpenStudy (dumbcow):

secA = 1/cosA = hypotenuse/adjacent = 13/12

OpenStudy (nice):

12 came from Pythagoras : (AC)^2 + (BC)^2 = (AB)^2

OpenStudy (anonymous):

Yay! Thanks(: